Because of an oil spill there is an oil sheen on the surface of the water in a lake. If the critical angle for a light ray leaving the water is 72, at what angle above the horizontal is the sun if sunlight reflected from the oily surface is completely polarized?
Answer #1

We need to find the Brewster angle first

\tan\theta_b=\frac{n_2}{n_1}

where \theta_b,n_2,n_1 is Brewster angle,index of refraction of denser medium,index of refraction of rarer medium

From Snell's law

n_1\sin\theta_1=n_2\sin\theta_2

At critical angle

\theta_2=90^\circ and \theta_1=72^\circ

\frac{n_2}{n_1}=0.95

Putting in the expression for Brewster angle

\tan\theta_b=0.95

\theta_b=43.56

Angle above the horizontal where sun is

90-43.56=46.44^\circ